# Generalization capabilities and robustness of hybrid machine learning models grounded in flow physics compared to purely deep learning models

### v2

@@ Abstract
This study  [^1] The following article has been submitted to Physics of Fluids. After it is published, it will be found at https://pubs.aip.org/aip/pof. investigates the generalization capabilities and robustness of purely deep learning (DL) models and hybrid models based on physical principles in fluid dynamics applications, specifically focusing on iteratively forecasting the temporal evolution of flow dynamics. Three autoregressive models were compared: a convolutional autoencoder combined with a convolutional LSTM (ConvLSTM), a variational autoencoder (VAE) combined with a ConvLSTM and a hybrid model that combines proper orthogonal decomposition (POD) with a LSTM (POD-DL). These models were tested on two high-dimensional, nonlinear datasets representing the velocity field of flow past a circular cylinder in both laminar and turbulent regimes. The study used latent dimension methods, enabling a bijective reduction of high-dimensional dynamics into a lower-order space to facilitate future predictions. While the VAE and ConvLSTM models accurately predicted laminar flow, the hybrid POD-DL model outperformed the others across both laminar and turbulent flow regimes. This success is attributed to the model’s ability to incorporate modal decomposition, reducing the dimensionality of the data, by a non-parametric method, and simplifying the forecasting component. By leveraging POD, the model not only gained insight into the underlying physics, improving prediction accuracy with less training data, but also reduce the number of trainable parameters as POD is non-parametric. The findings emphasize the potential of hybrid models, particularly those integrating modal decomposition and deep learning, in predicting complex flow dynamics.
@@ I Introduction
Fluid dynamics is a field with extensive applications across both industrial and natural processes where flows are involved. These applications range from combustion systems Le Clainche Martínez, Benocci, and Parente (2014), power plants, and airplanes Marusic et al. (2003), to the propulsion mechanisms of marine animals DeMont and Gosline (1988), blood flow in tumors Abadía-Heredia et al. (2021), and the wake of flying insects Muijres et al. (2005), among others. Given its broad applicability, there is significant interest in predicting these types of dynamics. However, they are governed by non-linear, high-dimensional systems of equations, making their resolution a challenging problem. In most situations, an analytical solution is not feasible, and the numerical resolution requires a discretization of the equations with a high number of degrees of freedom.
@@
This complexity is exacerbated in complex flows, namely flows in transitional and turbulent regime, which describe the majority of flows encountered in both industry and nature. These complex flows are characterized by the existence of multiple spatio-temporal scales within their dynamics. The number of scales increases exponentially with the Reynolds number (Re), a non-dimensional number comparing viscous and convective terms. In computational terms the cost of performing a numerical simulation, of all the spatio-temporal scales inside the flow, scales as $\hbox{Re}^{3}$,Karniadakis and Orszag (1993) which makes the numerical resolution of all the scales unfeasible for certain problems.
@@
Therefore, there is a high interest in reducing the computational cost associated with obtaining high-fidelity predictions of flow dynamics Brunton, Noack, and Koumoutsakos (2020). To address this issue, reduced order models (ROMs) have been proposed as an effective solution Rowley, Colonius, and Murray (2004). Where, they are designed to capture the most essential dynamics of a flow within a given context. There are two primary types of ROMs: intrusive projection-based ROMs and non-intrusive data-driven ROMs. Intrusive projection-based techniques, such as the Galerkin projection, aim to reduce the degrees of freedom in the governing equations by expanding them within a transformed space. This space is typically constructed using an orthogonal basis that retains only the most significant or energetic modes. The orthogonal basis is commonly derived through proper orthogonal decomposition (POD) Aubry (1991), a method first applied to analyze a turbulent flow by Lumley Lumley (1967). By projecting the governing equations onto this energetic basis, the ROM effectively captures the primary flow dynamics. The basis must be carefully computed depending on the specific problem Borrelli et al. (2022). Conversely, a non-intrusive data-driven ROM approach can infer the underlying physics from the data itself without requiring explicit knowledge of the governing equations. In this context, techniques such as POD Aubry (1991), spectral POD Sieber, Paschereit, and Oberleithner (2016), dynamic mode decomposition (DMD) Schmid (2010) and higher order dynamic mode decomposition (HODMD) Le Clainche and Vega (2017) are commonly utilized. A complete review of non-intrusive data-driven ROM techniques applied to fluid dynamics can be found in Begiashvili et al. (2023).
@@
The DL-ROM methodologies discussed above predict flow dynamics for new cases similar to those on which the model was trained, e.g., different boundary conditions or changes in parameters within the equations, such as the Reynolds number. However, in other scenarios, we might aim to forecast the flow dynamics over time based on a historical record of the dynamics, allowing a significant reduction in the computational cost associated with obtaining solutions required. This is the scope of this work.
Time series forecasting, applied to fluid dynamics, has been extensively studied in the literature, and different deep learning architectures and methodologies have been proposed. For example, Pathak et al. Pathak et al. (2018) used an echo-state-network based on recurrent neural networks to forecast the state evolution of a chaotic system. In another work, Vlachas et al. Vlachas et al. (2018) proposed the use of long short-term memory (LSTM) networks Hochreiter and Schmidhuber (1997) to forecast chaotic systems. In Han et al. (2019); Hasegawa et al. (2020), a DL architecture combining a convolutional autoencoder with LSTM networks was introduced to capture the spatial and temporal features of a spatio-temporal flow dynamics. Similarly, Nakamura et al. (2021) extended this approach to three-dimensional flow fields, enhancing the model’s ability to handle more complex spatial patterns.
@@
The use of autoencoders can be substitute by non-intrusive ROMs. For example, Ali et al. Ali et al. (2021) proposed a hybrid model combining dynamic mode decomposition with control (DMDc) and deep learning architectures. POD can also be employed to reduce the dimensionality, with forecasting performed on the POD coefficients to predict the temporal evolution of the POD modes. Several authors have studied these types of hybrid models, combining POD with deep learning architectures Mohan and Gaitonde (2018); Pawar et al. (2019); Abadía-Heredia et al. (2022); Xu et al. (2022); Corrochano et al. (2023). Moreover, a probabilistic forecasting version of these hybrid POD-DL models has been proposed by Maulik et al. (2020).
@@
This work compares three autoregressive methodologies to forecast flow dynamics from a small dataset, reducing the computational cost associated with computing the actual solution by numerical solvers. The dimensionality of this dataset is either reduced by POD, which identifies a series of coherent structures by means of modal decomposition, or autoencoders, which utilizes deep learning architectures to reduce the dimension.The compression capacity of autoencoders (non-linear) is much higher than the POD, however, they lack interpretability of the system physics Eivazi et al. (2022), except when using linear autoencoders, whose solution is similar to that of the POD modes Milano and Koumoutsakos (2002); Fukami, Fukagata, and Taira (2020); Fukami, Nakamura, and Fukagata (2020); Murata, Fukami, and Fukagata (2020); Muñoz et al. (2023). To the best of the authors’ knowledge, no comprehensive comparison of these existing methods has been conducted to date. This paper presents, for the first time, an in-depth evaluation of the capabilities of this class of models and demonstrates the advantages of generating hybrid models based on physics-informed approaches, like the POD-DL. The results indicate that the predictive power, generalization ability, and robustness of these hybrid models surpass those of purely deep-learning-based methods. This work represents a significant breakthrough in the field, achieving an optimal balance between computational efficiency—enabled by dimensionality reduction—and the generation of new, high-fidelity data.
@@ II Methodology
Where $J_{1}$ represents the components dimension of the dataset, as explained in Appendix A, and $J_{2}$, $J_{3}$, $J_{4}$ represent the spatial dimension in $x$-, $y$- and $z$-axis, respectively. In the context of point forecasting, the function $f$ represents the forecasting process, as follows,
@@
The probabilistic forecasting framework is similar to point forecasting, with the key difference being that instead of approximating a function, we aim to approximate the probability distribution $\mathbb{F}_{\bm{v}}$ underlying the forecasting process. To achieve this, a variational autoencoder (VAE) (Kingma and Welling, 2022) is employed, utilizing the Evidence Lower Bound (ELBO), defined in (5), as the loss function (Blei, Kucukelbir, and McAuliffe, 2017). In the following section we explain the proposed models in more detail, starting with the two models within the point forecast framework.
@@ II.1 Point forecasting: Residual autoencoder and POD-DL
This subsection presents the models that belong to the point forecasting methodology, where the objective is to approximate a function as described in eq. (1). The models presented reduce data dimensionality by means of an encoder (residual autoencoder) or POD (POD-DL). In the reduced-dimensional space, deep learning architectures are used to forecast the evolution of the flow. Finally, to recover the original (high-dimensional) space, the residual autoencoder uses a decoder and the POD-DL reverses the POD.
More precisely, in the POD-DL model, POD modes are calculated using singular value decomposition (SVD) Sirovich (1987), where only the most energetic POD modes representing the flow dynamics are retained. The predictions are then performed in the reduced basis of modes: the evolution of the POD temporal coefficients is calculated using neural networks. Here is where resides the principal advantage of the POD-DL model, in its substantial reduction in dimensionality, as it replaces high-dimensional snapshots with a simple vector representing the number of retained modes. This reduction is done by the SVD decomposition of the dataset into three matrices ($\bm{U}$, $\bm{\Sigma}$, $\bm{V}^{T}$), where $\bm{\Sigma}$ and $\bm{V}^{T}$ are used for forecasting, and $\bm{U}$ is kept as is to reconstruct the snapshots.
@@
The detailed architecture of the model, when applied to the laminar flow, is provided in Table 5. This architecture was already used in previous works Abadía-Heredia et al. (2022); Corrochano et al. (2023) in different applications to the ones presented in this article, illustrating the generalization capabilities of this model and the robustness of the architecture presented. Figure 1 provides a visual depiction of the POD-DL methodology.
@@
Figure 1: The POD-DL methodology can be represented as an encoder-decoder structure. However, in this case the encoder is just the matrix decomposition ($\bm{U},\bm{\Sigma},\bm{V}^{T}$) from SVD. The temporal information is formed by a matrix multiplication $\bm{M}_{t-n+1}^{t}=\bm{\Sigma}*(\bm{V}_{t-n+1}^{t})^{T}$, which is send to a DL model, usually a long-short term memory (LSTM), to predict the time-ahead sample, i.e., $\bm{M}_{t+1}$. Finally, the decoder reconstructs the snapshot by multiplying matrices $\bm{x}_{t+1}=\bm{U}*\bm{M}_{t+1}$.
@@
The second model, which is based on the works presented by Han et al. (2019); Hasegawa et al. (2020); Nakamura et al. (2021), uses a residual convolutional autoencoder, as illustrated in Figure 2. The detailed architecture of the model, when applied to the laminar flow, is provided in Tables 6, 7, and 8 in Appendix B.
@@
The encoder, on the one hand, consists of residual convolutional networks (ResCNN) (He et al., 2016), where the input and output of a CNN are added together. These ResCNNs are primarily defined by the identity and convolutional blocks, depicted in Figure 3. Both blocks add the input and output values, but with a key difference: the identity block neither reduces the size nor changes the number of channels of the input snapshot, while the convolutional block does. This architecture accelerates training and enhances robustness. The underlying principle is that it is easier for a block to represent the identity function (requiring only the convolutional kernels to be zero) than for the CNN to do so. On the other hand, the decoder is composed of transpose convolutional networks, which perform the inverse operation of convolutional kernels, allowing to go back from the low-dimensional latent space to the high-dimensional dynamics, as shown in Figure 2.
Figure 2: In this case both the encoder and decoder are composed by convolutional neural networks, specifically the encoder is composed by residual convolutional networks and the decoder is composed by transpose convolutional networks. Both of them catch the spatial correlation in data, while the hidden layer is formed by a ConvLSTM network that captures the temporal correlation in the low-dimensional latent space.
@@
Note that both the encoder and decoder process only the spatial dimensions of the snapshots and do not alter the temporal dimension. In Appendix B, Figure 21 displays some outputs from the encoder when incoming snapshots are from the laminar flow. Unlike the latent variables obtained from SVD, these variables are not hierarchically ordered based on the energy they contain. However, several methods have been proposed to sort the latent variables produced by an autoencoder (Fukami, Nakamura, and Fukagata, 2020; Muñoz et al., 2023), providing information on which latent variable contributes most to the overall dynamics.
@@
A hidden layer composed of a convolutional LSTM layer (ConvLSTM) (Shi et al., 2015) is used to capture the temporal correlation in the input sequence of snapshots $\{\bm{v}_{i-n+1},\bm{v}_{i-n+2},\dots,\bm{v}_{i}\}$. For detailed architectural information, refer to Table 7. Note in Figure 2, this hidden layer is situated between the encoder and decoder to forecast on the latent variables.

### v4

# Generalization capabilities and robustness of hybrid models grounded in physics compared to purely deep learning models

@@ Abstract
This study  [^1] This article has been submitted to Physics of Fluid. After it is published, it will be found at https://pubs.aip.org/aip/pof. investigates the generalization capabilities and robustness of purely deep learning (DL) models and hybrid models based on physical principles in fluid dynamics applications, specifically focusing on iteratively forecasting the temporal evolution of flow dynamics. Three autoregressive models were compared: a hybrid model (POD-DL) that combines proper orthogonal decomposition (POD) with a long-short term memory (LSTM) layer, a convolutional autoencoder combined with a convolutional LSTM (ConvLSTM) layer and a variational autoencoder (VAE) combined with a ConvLSTM layer. These models were tested on two high-dimensional, nonlinear datasets representing the velocity field of flow past a circular cylinder in both laminar and turbulent regimes. The study used latent dimension methods, enabling a bijective reduction of high-dimensional dynamics into a lower-order space to facilitate future predictions. While the VAE and ConvLSTM models accurately predicted laminar flow, the hybrid POD-DL model outperformed the others across both laminar and turbulent flow regimes. This success is attributed to the model’s ability to incorporate modal decomposition, reducing the dimensionality of the data, by a non-parametric method, and simplifying the forecasting component. By leveraging POD, the model not only gained insight into the underlying physics, improving prediction accuracy with less training data, but also reduce the number of trainable parameters as POD is non-parametric. The findings emphasize the potential of hybrid models, particularly those integrating modal decomposition and deep learning, in predicting complex flow dynamics.
@@ I Introduction
Fluid dynamics is the study of fluid motion and is fundamental to various natural and industrial processes involving flow, including combustion systems Le Clainche Martínez, Benocci, and Parente (2014), power generation, aerodynamics Marusic et al. (2003), marine propulsion DeMont and Gosline (1988), blood flow in tumors Abadía-Heredia et al. (2021), and the wakes of flying insects Muijres et al. (2005), among others. This broad range of applications has driven significant interest in the characterization of fluid flows. However, they are governed by nonlinear, high-dimensional system of equations, posing substantial challenges for their resolution. In most cases, analytical solutions are intractable, and numerical approaches require discretization schemes with a large number of degrees of freedom, leading to high computational costs.
@@
This complexity is exacerbated in complex flows, namely flows in transitional and turbulent regime, which describe the majority of flows encountered in both industry and nature. These complex flows are characterized by the existence of multiple spatio-temporal scales within their dynamics. The number of scales increases nearly quadratic with the Reynolds number (Re), a non-dimensional number defined as $\hbox{Re}=VD/\nu$, where $V$ and $D$ are the velocity and the characteristic length of the problem, respectively, and $\nu$ denotes the kinematic viscosity. However, resolving all flow scales numerically can become computationally prohibitive for many problems.
@@
Therefore, there is a high interest in reducing the computational cost associated with obtaining high-fidelity predictions of flow dynamics Brunton, Noack, and Koumoutsakos (2020). To address this issue, reduced order models (ROMs) have been proposed as an effective solution Rowley, Colonius, and Murray (2004). They are designed to capture the most essential dynamics of a flow within a given context. There are two primary types of ROMs: intrusive projection-based ROMs and non-intrusive data-driven ROMs. Intrusive projection-based techniques, such as the Galerkin projection, aim to reduce the degrees of freedom in the governing equations by expanding them within a transformed space. This space is typically constructed using an orthogonal basis that retains only the most significant or energetic modes. The orthogonal basis is commonly derived through proper orthogonal decomposition (POD) Aubry (1991), a method first applied to analyze a turbulent flow by Lumley Lumley (1967). By projecting the governing equations onto this energetic basis, the ROM effectively captures the primary flow dynamics. The basis must be carefully computed depending on the specific problem Borrelli et al. (2022). Conversely, a non-intrusive data-driven ROM approach can infer the underlying physics from the data itself without requiring explicit knowledge of the governing equations. In this context, techniques such as POD Aubry (1991), spectral POD Sieber, Paschereit, and Oberleithner (2016), dynamic mode decomposition (DMD) Schmid (2010) and higher order dynamic mode decomposition (HODMD) Le Clainche and Vega (2017) are commonly utilized. A complete review of non-intrusive data-driven ROM techniques applied to fluid dynamics can be found in Begiashvili et al. (2023).
@@
In other scenarios these DL-ROM methodologies can be used to forecast the flow dynamics based on a historical record, allowing a significant reduction in the computational cost associated with obtaining the solutions required. This is the scope of this work.
Time series forecasting, applied to fluid dynamics, has been extensively studied in the literature, and different deep learning architectures and methodologies have been proposed. For example, Pathak et al. Pathak et al. (2018) used an echo-state-network based on recurrent neural networks to forecast the state evolution of a chaotic system. In another work, Vlachas et al. Vlachas et al. (2018) proposed the use of long short-term memory (LSTM) networks Hochreiter and Schmidhuber (1997) to forecast chaotic systems. In works proposed by R. Han et al. Han et al. (2019) and K. Hasegawa et al. Hasegawa et al. (2020), a DL architecture combining a convolutional autoencoder with LSTM networks was introduced to capture the spatial and temporal features of a spatio-temporal flow dynamics. Similarly, T. Nakamura et al. Nakamura et al. (2021) extended this approach to three-dimensional flow fields, enhancing the model’s ability to handle more complex spatial patterns.
@@
The use of DL autoencoders can be substitute by non-intrusive ROMs. For example, Ali et al. Ali et al. (2021) proposed a hybrid model combining dynamic mode decomposition with control (DMDc) and deep learning architectures. POD can also be employed to reduce the dimensionality, with forecasting performed on the POD coefficients to predict the temporal evolution of the POD modes. Several authors have studied these types of hybrid models, combining POD with deep learning architectures Mohan and Gaitonde (2018); Pawar et al. (2019); Parish and Carlberg (2020); Abadía-Heredia et al. (2022); Xu et al. (2022); Corrochano et al. (2023). Moreover, a probabilistic forecasting version of these hybrid POD-DL models has been proposed by Maulik et al. (2020).
@@
This study evaluates three autoregressive methodologies for forecasting spatiotemporal flow dynamics using a small dataset, with the aim of reducing computational costs compared to numerically solving the governing equations. The selected models are grounded in the contemporary literature about spatiotemporal forecasting and include the following: (1) A residual autoencoder, chosen for its proven efficacy in handling spatiotemporal data by integrating a convolutional autoencoder for spatial feature extraction with a convolutional long short-term memory (ConvLSTM) network for temporal forecasting; (2) a variational autoencoder (VAE), which adopts a similar architecture—combining a convolutional autoencoder with an LSTM network—while leveraging a probabilistic framework; and (3) a POD-DL model, which integrates modal decomposition techniques with deep learning architectures to incorporate physical insights and enhance predictive accuracy.
Another objective of this study is to compare various methods of dimensionality reduction for the creation of latent dimensions. Specifically, the performance of deep learning autoencoders—encompassing both the residual and the variational autoencoders—is evaluated against POD, which is employed by the POD-DL model.
Deep learning autoencoders rely entirely on neural networks to compress high-dimensional data into a lower-dimensional latent space, enabling the capture of nonlinear relationships within the spatial domain. In contrast, POD achieves dimensionality reduction by projecting data onto an orthogonal basis constructed from a set of POD modes, which effectively represent the most energetically significant structures in the system’s dynamics.
While nonlinear DL autoencoders generally exhibit superior compression capabilities compared to POD, they often lack interpretability regarding the underlying physical phenomena Eivazi et al. (2022); Fukami, Fukagata, and Taira (2020); Fukami, Nakamura, and Fukagata (2020); Murata, Fukami, and Fukagata (2020); Muñoz et al. (2023). An exception to this is the linear autoencoder, whose solutions have been observed to closely resemble those of POD modes Milano and Koumoutsakos (2002). To maintain clarity, the term "autoencoder" will hereafter refer exclusively to DL autoencoders, unless explicitly stated otherwise.
To the best of the authors’ knowledge, no comprehensive comparison of these existing methods has been conducted to date. This paper presents, for the first time, an in-depth evaluation of the capabilities of this class of models and demonstrates the advantages of generating hybrid models based on physics-informed approaches, like the POD-DL. The results indicate that the predictive power, generalization ability, and robustness of these hybrid models surpass those of purely deep-learning-based methods. This work represents a significant breakthrough in the field, achieving an optimal balance between computational efficiency—enabled by dimensionality reduction—and the generation of new, high-fidelity data.
@@ II Methodology
Here, $J_{1}$ represents the components dimension of the dataset (i.e., different velocity components, different Reynolds numbers, different initial conditions, etc.). The indices $J_{2}$, $J_{3}$, $J_{4}$ correspond to the spatial discretization along the $x$-, $y$- and $z$-axis, respectively. The variable $q$ denotes the length of the historical data used for prediction, while the value $1$ indicates that the objective is to forecast a single time-step ahead. The computation of more than one prediction, as shown in Sec. III, is obtained by an autoregressive procedure. A more detailed explanation of the dataset structure and preprocessing steps is provided in Appendix A.
More precisely this function $f$ represents the forecasting process, as follows,
@@
The probabilistic forecasting framework is similar to point forecasting, with the key difference being that instead of approximating a function, we aim to approximate the probability distribution $\mathbb{F}_{\bm{v}}$ underlying the forecasting process. To achieve this, a variational autoencoder (VAE) (Kingma and Welling, 2022) is employed. In the following section we explain the proposed models in more detail, starting with the two models within the point forecast framework.
@@ II.1 Point forecasting: Residual autoencoder and POD-DL
This subsection presents the models that belong to the point forecasting methodology, where the objective is to do a deterministic prediction by approximating a function as described in eq. (1). The models presented reduce data dimensionality by means of an encoder (residual autoencoder) or POD (POD-DL). In the reduced-dimensional space, deep learning architectures are used to forecast the evolution of the flow. Finally, to recover the original (high-dimensional) space, the residual autoencoder uses a decoder and the POD-DL reverses the POD.
More precisely, in the POD-DL model, singular value decomposition (SVD) Sirovich (1987) is applied to decompose a high-dimensional dataset $\bm{D}$ in a product of matrices, as follows,
$\bm{D}=\bm{U}\bm{\Sigma}\bm{V}^{T}.$ (4)
Here, the matrix $\bm{U}$ represents the POD modes, the diagonal matrix $\bm{\Sigma}$ contains the singular values, and $\bm{V}^{T}$ corresponds to the POD coefficients.
In most applications is enough to retain only the most energetic POD modes (i.e., columns of matrix $\bm{U}$) representing important structures (i.e., coherent structures) within the flow dynamics. This truncation of modes is usually done by looking at the singular values (i.e., diagonal elements of matrix $\bm{\Sigma}$) that contain the relevance of each POD mode within the flow dynamics. The temporal evolution of these POD modes is carried out by the POD coefficients (i.e., columns of matrix $\bm{V}^{T}$).
The forecasting is then performed in the weighted POD coefficients, $\bm{\Sigma}\bm{V}^{T}$, where their temporal evolution is predicted using deep learning architectures, drawing parallels to Galerkin methods.
Here is where resides the principal advantages of the POD-DL model in contrast to the other models: (1) truncating the POD modes enables forecasting to focus specifically on the most important flow scales for a required problem, as in most cases there is no need to solve the smaller flow scales, generally related to uncorrelated flow dynamics; (2) SVD is a computationally efficient approach to compute the POD modes; and (3) by going from a high-dimensional tensor to a matrix (i.e., $\bm{\Sigma}\bm{V}^{T}$) where columns represent the temporal dimension, POD simplifies the model architecture required for forecasting, reducing the risk of overfitting and improving generalization.
In this work the deep learning architecture used to predict the evolution of the POD temporal coefficients is a long-short term memory (LSTM), where in a previous study Parish and Carlberg (2020) was observed to outperform traditional forecasting methods in both accuracy and computational efficiency. This advantage arises primarily because traditional methods often require significant computational resources and struggle to generalize to new or unseen conditions. This was also observed in the M4 forecasting competition where machine learning models outperformed the classical probabilistic models Makridakis, Spiliotis, and Assimakopoulos (2018).
@@
The detailed architecture of the POD-DL model is provided in Table 6. This architecture was already used in previous works Parish and Carlberg (2020); Abadía-Heredia et al. (2022); Corrochano et al. (2023) in different applications to the ones presented in this article, illustrating the generalization capabilities of this model and the robustness of the architecture presented. Figure 1 provides a visual depiction of the POD-DL methodology.
@@
Figure 1: Overview of the architecture of the POD-DL model. Note, it can be represented as an encoder-decoder structure. However, in this case the encoder is just the matrix decomposition ($\bm{U},\bm{\Sigma},\bm{V}^{T}$) from SVD. The temporal information is formed by a matrix multiplication $\bm{M}_{t-n+1}^{t}=\bm{\Sigma}*(\bm{V}_{t-n+1}^{t})^{T}$, which is send to a DL model, usually a long-short term memory (LSTM), to predict the time-ahead sample, i.e., $\bm{M}_{t+1}$. Finally, the decoder reconstructs the snapshot by multiplying matrices $\bm{x}_{t+1}=\bm{U}*\bm{M}_{t+1}$.
@@
The second model, which is based on the works presented by Han et al. (2019); Hasegawa et al. (2020); Nakamura et al. (2021), uses a residual convolutional autoencoder, as illustrated in Fig. 2. The detailed architecture of the model, when applied to the laminar flow, is provided in Tables 7, 8, and 9 in Appendix D.
@@
The encoder, on the one hand, consists of residual convolutional networks (ResCNN) (He et al., 2016), where the input and output of a CNN are added together. These ResCNNs are primarily defined by the identity and convolutional blocks, depicted in Fig. 4. Both blocks add the input and output values, but with a key difference: the identity block neither reduces the size nor changes the number of channels of the input snapshot, while the convolutional block does. This architecture accelerates training and enhances robustness. The underlying principle is that it is easier for a block to represent the identity function (requiring only the convolutional kernels to be zero) than for the CNN to do so. On the other hand, the decoder is composed of transpose convolutional networks, which perform the inverse operation of convolutional kernels, allowing to go back from the low-dimensional latent space to the high-dimensional dynamics, as shown in Figure 2.
Figure 2: Overview of the architecture of the residual autoencoder. Unlike the POD-DL, the encoder and decoder of this model are composed of convolutional neural networks.
@@
Note that both the encoder and decoder process only the spatial dimensions of the snapshots and do not alter the temporal dimension. In Appendix D, Fig. 26 displays some outputs from the encoder when incoming snapshots are from the laminar flow. Unlike the latent variables obtained from SVD, these variables are not hierarchically ordered based on the energy they contain. However, several methods have been proposed to sort the latent variables produced by an autoencoder (Fukami, Nakamura, and Fukagata, 2020; Muñoz et al., 2023), providing information on which latent variable contributes most to the overall dynamics.
@@
A hidden layer composed of a convolutional LSTM layer (ConvLSTM) (Shi et al., 2015) is used to capture the temporal correlation in the input sequence of snapshots $\{\bm{v}_{i-n+1},\bm{v}_{i-n+2},\dots,\bm{v}_{i}\}$. For detailed architectural information, refer to Table 8. Note in Fig. 2, this hidden layer is situated between the encoder and decoder to forecast on the latent variables.